A spark plug manufacturing company produces a type A part using a Kanban system. The demand for the type A part is 20 /hour and the size of the container used is 5. If the lead time is uniformly distributed with [2,12], how many Kanbans are required for Type A part to ensure that parts are available for at least 95% of the time they are requested?
Answer #1

Lead time = 2 + 0.95*(12 - 2) = 11.5

For given 95% service level = 20*11.5 = 230 units

Given the size of the container used is 5, therefore the number of kanbans required to cover the lead time demand = 230 / 5 = 46